# Partial Rewriting and Value Interpretation of Logically Constrained Terms

### v1

@@ 6 Characterization of Partial and Most General Constrained Reductions by Value Instantiation
Let $\rho:\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ be a left-linear and left-value-free constrained rewrite rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$. Suppose that for any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ there exists $v$ such that $\langle u|_{p},v|_{p}\rangle\in\{\langle\ell\sigma,r\sigma\rangle\mid Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma),\sigma\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\}$, where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$.
@@
It remains to show (iv). For this, suppose $\xi$ is a valuation such that $\vDash_{\mathcal{M},\xi}\exists\vec{x}.\ \varphi$. Take a substitution $\hat{\xi}=\xi|_{X}$. Then $\hat{\xi}$ is $X$-valued and $\hat{\xi}\vDash\exists\vec{x}.\ \varphi$ by $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X$. Also, as $\{\vec{x}\}\cap X=\varnothing$, we have ${\exists\vec{x}.\ \varphi\hat{\xi}}]\!]_{\mathcal{M}}=\mathsf{true}$. Let $u^{\prime}=s\hat{\xi}$. Then, $u^{\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$. Thus, by our assumption, there exists a substitution $\sigma^{\prime}$ such that (1’) $u^{\prime}|_{p}=\ell\sigma^{\prime}$, (2’) $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma^{\prime})$, and (3’) $\sigma^{\prime}\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$.
@@
Let $x\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)$, i.e., $x\in\mathcal{V}\mathsf{ar}(\pi\gamma)\setminus(\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell))$. Then $x\in\{\gamma(z)\in\mathcal{V}\mid z\in\mathcal{V}\mathsf{ar}(\ell)\cap Z\}$. Thus, let $q,z$ be such that $z\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, $x=\gamma(z)$, and $\ell|_{q}=z$. Then $\sigma^{\prime}(z)\in\mathcal{V}\mathsf{al}$ by (2’), and hence using (1’), we have $u^{\prime}|_{q}=\ell\sigma^{\prime}|_{q}=(\ell|_{q})\sigma^{\prime}=\sigma^{\prime}(z)\in\mathcal{V}\mathsf{al}$. Then, as $s\hat{\xi}=u^{\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$, we have $s|_{q}\in X\cup\mathcal{V}\mathsf{al}$. Then, $\hat{\xi}(x)=\hat{\xi}(\gamma(z))=\ell|_{q}\gamma\hat{\xi}=(\ell\gamma|_{q})\hat{\xi}=(s|_{q})\hat{\xi}\in\mathcal{V}\mathsf{al}$, as $\hat{\xi}=\xi|_{X}$. Thus, for any $x\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)$, we have $\hat{\xi}(x)\in\mathcal{V}\mathsf{al}$. Furthermore, $\hat{\xi}(\gamma(z))=\hat{\xi}(x)=(s|_{q})\hat{\xi}=u^{\prime}|_{q}=\sigma^{\prime}(z)$. Thus, $\vDash_{\mathcal{M},\hat{\xi}}\exists\vec{z}.\ \pi\gamma$ by (3’). Hence, $\vDash_{\mathcal{M},\xi}\exists\vec{z}.\ \pi\gamma$. Thus, it follows that $\vDash_{\mathcal{M},\xi}(\exists\vec{x}.\ \varphi)\Rightarrow(\exists\vec{z}.\ \pi\gamma)$ for any valuation $\xi$, and therefore (iv) holds.
Now it follows that Lemma˜4.12 can be strengthened with the converse direction by additional condition that the rewrite rule is left-value-free.
Corollary 6.10.
  1. 1.
  2. 2.
Proof 6.11.
(2 $\Longrightarrow$ 1) follows by Lemma˜4.12. (1 $\Longrightarrow$ 2) follows from Lemma˜6.8 and Lemma˜5.3 3.
Now we arrive at a characterization of normal forms with respect to most general constrained reduction in terms of value interpretation. This is similar to the notion of normal form of constrained terms in [15, Definition 1].
Theorem 6.12 (Normal Forms of Most General Constrained Reduction).
Proof 6.13.
[…]
@@
($\Longrightarrow$) We prove the contraposition. Suppose there exist $\rho\in\mathcal{R}$ and $p\in\mathcal{P}\mathsf{os}(s)$ such that for any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$ for some $v$. Then by Lemma˜6.8, $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\to^{p}_{\rho}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ for some $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$.
Note that the theorem above can be reformulated as follows. The following statements are equivalent: (i) $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\to_{\mathcal{R}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$ for some $\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$, and (ii) there exists $\rho\in\mathcal{R}$ and position $p\in\mathcal{P}\mathsf{os}(s)$ such that for any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ there exists $v$ such that $u\to^{p}_{\langle\!\langle\rho\rangle\!\rangle}v$. In contrast to the characterization of partial constrained reduction (Theorem˜6.5), our characterization of most general constrained reduction refers to reducibility of all $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ with respect to a fixed position and a fixed rewrite rule.
Example 6.14.
Let us recall the LCTRS $\mathcal{R}$ in Example˜1.1:
[…]
Consider the rewrite steps starting from the constrained term $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\hskip 0.85358pt]$. Note that we have $\langle\!\langle\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}=\{\mathsf{sum}(0),\ldots,\mathsf{sum}(4)\}$, $\langle\!\langle\rho_{1}\rangle\!\rangle=\{\mathsf{sum}(n)\to 0\mid 0\geqslant n\in\mathbb{Z}\}$, and $\langle\!\langle\rho_{2}\rangle\!\rangle=\{\mathsf{sum}(n)\to n+\mathsf{sum}(n+-1)\mid 0<n\in\mathbb{Z}\}$. We further have $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\hskip 0.85358pt]\leadsto^{\epsilon}_{\rho_{1}}\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ 0\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\land 0\geqslant x\hskip 0.85358pt]$, and $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\hskip 0.85358pt]\leadsto^{\epsilon}_{\rho_{2}}\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ x+\mathsf{sum}(x+-1)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\land 0<x\hskip 0.85358pt]$. By Theorem˜6.5, the former is characterized by $\mathsf{sum}(0)\to^{\epsilon}_{\langle\!\langle\rho\rangle\!\rangle}0$ and the latter by $\mathsf{sum}(4)\to^{\epsilon}_{\langle\!\langle\rho\rangle\!\rangle}4+\mathsf{sum}(4+-1)$. Also none of the TRSs $\langle\!\langle\rho_{1}\rangle\!\rangle,\langle\!\langle\rho_{2}\rangle\!\rangle$ can reduce all of the terms $\mathsf{sum}(0),\ldots,\mathsf{sum}(4)$. By Theorem˜6.12, we obtain that $\Pi\{\hskip 0.85358ptx\hskip 0.85358pt\}.\ \mathsf{sum}(x)\penalty 10000\ [\hskip 0.85358pt0\leqslant x\land x\leqslant 4\hskip 0.85358pt]$ is a normal form with respect to $\to_{\mathcal{R}}$.
7 Related Work
Partial constrained reduction implicitly appears in various forms in the applications and the analyses of LCTRSs and of a related formalism of rewriting with built-in data structures.
Critical Pair Analysis
Critical pair analysis is a standard technique for checking confluence of LCTRSs. Many approaches rely on closure conditions for constrained critical pairs. Splitting a constrained critical pair [15] partitions its set of instances into subsets, each of which may admit a rewrite step. The goal is that, after splitting, a concrete rewrite step becomes applicable and satisfies the closing condition, which corresponds to a partial rewrite step on the original constrained critical pair. Partial rewrite steps can also be used to determine whether such a split is possible, as exploited to some extent in tools like crest [15].
Partial rewrite steps can also be used to detect non-confluence in LCTRSs, since it suffices that an instance of a critical pair diverges. In this case, explicit splitting is not required: a partial step could directly expose the subset of instances responsible for divergence. Tools like crest exploit this principle to efficiently find counterexamples without actual partitioning.
Constrained Rewriting Induction
An equation is said to be an inductive theorem of an equational theory when it is valid in the initial algebra of the theory; in other words, when all ground instantiations of the equation are equational consequences of the theory (e.g. [5]). Rewriting induction, that has been initiated in [12], is an approach to verify inductive theorems based on rewriting techniques. The notion of inductive theorems and the rewriting induction approach have been incorporated to the LCTRS formalism in [4, 10], where the latter has been called “constrained rewriting induction.” A key inference rule in rewriting induction is the rule Expand, which consists of a collection of narrowing steps applied at an appropriate position in an equation using all rules applicable at that position. A narrowing step is a rewrite step that may involve a minimal instantiation of the target term, replacing matching with unification. In constrained rewriting induction, instead of using only pure narrowing steps, a combination of narrowing steps and partial constrained rewrite steps is employed. This approach collects the reducts of all possible rewrite steps for minimal instantiations, while potentially adding constraints to the constrained equation at an appropriate position. In addition, although not covered in [4], partial constrained reduction may be useful to strengthen the Disprove rule for refuting a conjecture in constrained rewriting induction, similarly to the part on critical pair analysis above.
Constrained Narrowing
“Constrained Narrowing” [3, 6] narrows both, instances of non-logical variables and constraints for logical variables: To apply a constrained rewrite rule to a constrained term, non-logical variables of the constrained term are minimally instantiated, and the constraint of the rule is added to the constraint of the constrained term. Thus, constrained narrowing can be considered as “Narrowing” + “Partial Rewriting”.
Rewriting Modulo SMT
“Rewriting Modulo SMT” [13] is a sophisticated and intricate rewriting formalism (in a broad sense) that is based on a framework and terminology different from that of LCTRSs. Nevertheless, it shares key features with LCTRSs, such as the ability to handle built-in data structures and the use of constrained rewrite rules and rewriting on constrained terms. Although precise comparisons between the two formalisms are beyond the scope of this paper, roughly speaking, this formalism is designed to model state transition systems for more application-oriented specifications. In contrast to LCTRSs, it includes extended features such as order-sorted signatures and rewriting modulo equations, as well as restrictions like rewriting only at the root position, focusing on initial models, and a hierarchical signature for term sorts over theory sorts. We would like to highlight two noteworthy correspondences between results in [13] and ours. First, the basic rewrite relation (referred to as the symbolic rewrite relation in [13]) corresponds more closely to partial constrained rewriting rather than to most general constrained rewriting. Second, papers establish soundness and completeness results with respect to the standard interpretations. Readers may compare the soundness results (Lemma˜4.8 and Lemma˜6.1 versus Theorem 1 of [13]) and completeness results (Lemma˜6.3 and Example˜6.7 versus Theorem 2 of [13]). Overall, our results help to clarify the differences between the rewriting modulo SMT and LCTRS formalisms and highlight the respective advantages of each.
8 Conclusion
Appendix A Omitted Proofs

### v2

@@ Appendix A Omitted Proofs
Let $\rho:\Pi Z.\ \ell\rightarrow r\penalty 10000\ [\hskip 0.85358pt\pi\hskip 0.85358pt]$ be a left-linear and left-value-free constrained rewrite rule satisfying $\mathcal{V}\mathsf{ar}(\rho)\cap\mathcal{V}\mathsf{ar}(s,\varphi)=\varnothing$. Suppose that for any $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$ there exists $v$ such that $\langle u|_{p},v|_{p}\rangle\in\{\hskip 0.85358pt\langle\ell\sigma,r\sigma\rangle\mid Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma),\sigma\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi\hskip 0.85358pt\}$, where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$.
@@
It remains to show (iv). For this, suppose $\xi$ is a valuation such that $\vDash_{\mathcal{M},\xi}\exists\vec{x}.\ \varphi$. Take a substitution $\hat{\xi}=\xi|_{X}$. Then $\hat{\xi}$ is $X$-valued and $\hat{\xi}\vDash\exists\vec{x}.\ \varphi$ by $\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{x}.\ \varphi)\subseteq X$. Also, as $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}\cap X=\varnothing$, we have ${\exists\vec{x}.\ \varphi\hat{\xi}}]\!]_{\mathcal{M}}=\mathsf{true}$. Let $u^{\prime}=s\hat{\xi}$. Then, $u^{\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$. Thus, by our assumption, there exists a substitution $\sigma^{\prime}$ such that (1’) $u^{\prime}|_{p}=\ell\sigma^{\prime}$, (2’) $Z\cap\mathcal{V}\mathsf{ar}(\ell,r)\subseteq\mathcal{V}\mathcal{D}\mathsf{om}(\sigma^{\prime})$, and (3’) $\sigma^{\prime}\vDash_{\mathcal{M}}\exists\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}.\ \pi$ where $\{\hskip 0.85358pt\vec{z}\mkern 2.0mu\vphantom{z}^{\prime}\hskip 0.85358pt\}=\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell,r)$.
@@
Let $x\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)$, i.e., $x\in\mathcal{V}\mathsf{ar}(\pi\gamma)\setminus(\mathcal{V}\mathsf{ar}(\pi)\setminus\mathcal{V}\mathsf{ar}(\ell))$. Then $x\in\{\hskip 0.85358pt\gamma(z)\in\mathcal{V}\mid z\in\mathcal{V}\mathsf{ar}(\ell)\cap Z\hskip 0.85358pt\}$. Thus, let $q,z$ be such that $z\in\mathcal{V}\mathsf{ar}(\ell)\cap Z$, $x=\gamma(z)$, and $\ell|_{q}=z$. Then $\sigma^{\prime}(z)\in\mathcal{V}\mathsf{al}$ by (2’), and hence using (1’), we have $u^{\prime}|_{q}=\ell\sigma^{\prime}|_{q}=(\ell|_{q})\sigma^{\prime}=\sigma^{\prime}(z)\in\mathcal{V}\mathsf{al}$. Then, as $s\hat{\xi}=u^{\prime}\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle_{\mathsf{v}}$, we have $s|_{q}\in X\cup\mathcal{V}\mathsf{al}$. Then, $\hat{\xi}(x)=\hat{\xi}(\gamma(z))=\ell|_{q}\gamma\hat{\xi}=(\ell\gamma|_{q})\hat{\xi}=(s|_{q})\hat{\xi}\in\mathcal{V}\mathsf{al}$, as $\hat{\xi}=\xi|_{X}$. Thus, for any $x\in\mathcal{F}\mathcal{V}\mathsf{ar}(\exists\vec{z}.\ \pi\gamma)$, we have $\hat{\xi}(x)\in\mathcal{V}\mathsf{al}$. Furthermore, $\hat{\xi}(\gamma(z))=\hat{\xi}(x)=(s|_{q})\hat{\xi}=u^{\prime}|_{q}=\sigma^{\prime}(z)$. Thus, $\vDash_{\mathcal{M},\hat{\xi}}\exists\vec{z}.\ \pi\gamma$ by (3’). Hence, $\vDash_{\mathcal{M},\xi}\exists\vec{z}.\ \pi\gamma$. Thus, it follows that $\vDash_{\mathcal{M},\xi}(\exists\vec{x}.\ \varphi)\Rightarrow(\exists\vec{z}.\ \pi\gamma)$ for any valuation $\xi$, and therefore (iv) holds.
See 6.12
Proof A.14.
@@
We first note that for any non-constrained term $u$, $u\to_{\rho}v$ if and only if $u\to_{\langle\!\langle\rho\rangle\!\rangle}v$. We show a contraposition. Suppose $\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\leadsto_{\mathcal{R}}\Pi Y.\ t\penalty 10000\ [\hskip 0.85358pt\exists\vec{y}.\ \psi\hskip 0.85358pt]$. Then, by Lemma˜4.7, there exists $u\in\langle\!\langle\Pi X.\ s\penalty 10000\ [\hskip 0.85358pt\exists\vec{x}.\ \varphi\hskip 0.85358pt]\rangle\!\rangle$ such that $u\to_{\langle\!\langle\mathcal{R}\rangle\!\rangle}v$ for some $v$. Thus, $u\to_{\mathcal{R}}v$ for some $v$.